Two vehicles are approaching an intersection. One is a 2500-π‘˜π‘” pickup traveling at 14.0π‘š/𝑠 from east to west (the βˆ’π‘₯-direction), and the other is a 1500-π‘˜π‘” sedan going from south to north (the +𝑦-direction) at 23.0 π‘š/𝑠. (a) Find the π‘₯- and 𝑦-components of the net momentum of this system. (b) What are the magnitude and direction of the net momentum?


1
Expert's answer
2021-12-12T16:40:25-0500
"p_x=-(2500)(14)=-35000\frac{kgm}{s}\\
p_y=(1500)(23)=34500\frac{kgm}{s}\\
p=\sqrt{35000^2+34500^2}=49100\frac{kgm}{s}\\\theta=\arctan{\frac{345}{350}}=44.6\degree"